A portfolio consists of equity and bonds. The two assets are equally weighted (i.e. equal amounts are invested in each of the two assets). The risk of equity is 17.0%, while the risk of bonds is 8.1%. Given that the correlation between the two assets is 0.2, what is the risk of the portfolio (as a standard deviation)?
Answer #1

Equally weighted portfolio means that the weight of each asset is 50% or 0.5

Weight of equity in the portfolio = w1 = 0.5

Weight of bonds in the portfolio = w2 = 0.5

Risk or standard deviation of equity = σ1 = 17%

Risk or standard deviation of bonds = σ2 = 8.1%

Correlation between the two assets (equity and bond) = ρ = 0.2

Variance of the portfolio is calculated using the formula:

Variance of the portfolio = σP2 = w1212 + w2222 + 2*w1*w2*ρ*σ12

σP2 = 0.52*(17%)2 + 0.52*(8.1%)2 + 2*0.5*0.5*0.2*17%*8.1% = 0.007225+0.00164025+0.001377 = 0.01024225

Standard deviation or risk is square-root of the variance

Standard deviation or risk of the portfolio = σP = 0.010242251/2 = 0.101204001897158 = 10.1204001897158 %

Answer -> Risk of the portfolio = 10.1204001897158 %